Air is contained in a piston cylinder device that has a resistance heater inside. The atmosphere exerts a pressure of 100 kpa on the piston of which it has a mass of 45 kg and a transverse area of ​​0.9 m². An electric current passes through the resistance of the air volume slowly increases by 0.45 m³ while the pressure remains constant. The mass of the air is 0.27 kg and its specific internal energy increases 42 kJ / kg, the air and the piston are at rest in the initial state. The piston is a good ceramic insulator. Ignore the friction between the piston and the cylinder and consider the acceleration of gravity as 9.81 m/s2. Determine the heat transfer in KJ to the system consisting of:
Air only
Air and piston

- - - Air Hradece
Answer #1

first law t thrody namics AU= W+ A ५२ ४ AU= x0-27 K9 K3 11340 J XAV t 3 6-9m2 45000 J 220.725 J A8=-W+AU -45000 +220.725+ 113